### v1

# On Computability of Computable Problems

Abstract:Computational problems are classified into computable and uncomputable problems.If there exists an effective procedure (algorithm) to compute a problem then the problem is computable otherwise it is uncomputable.Turing machines can execute any algorithm therefore every computable problem is Turing computable.There are some variants of Turing machine that appear computationally more powerful but all these variants have been proven equally powerful.The main objective of this work is to revisit and examine the computational power of different variants of Turing machines at very fine-grain level.We achieve this objective by constructing a transform technique for Turing computable problems that transforms computable problems into another type of problems, and then we try to compute the transformed problems through different variants of Turing machine.This paper shows the existence of a realizable computational scheme that can establish a framework to analyze computational characteristics of different variants of Turing machine at infinitesimal scale.
Key Words: Turing machine, Stipulation machine, Computability, Alphabet, Tape
1 Introduction
Discovery of paradoxes during attempts to axiomatize Mathematics played important role in foundational crisis of Mathematics in late 19th and early 20th century.There were three major schools of thought to deal with this crisis, i.e., intuitionism, logisim, and formalism.Main opinions about this foundational crisis were turned into Mathematical research projects.David Hilbert, one of the most notable formalists, presented one such project that is known as Hilbert’s program [1].Entscheidungsproblem, a decision problem of first order logic, was posed by David Hilbert and Wilhelm Ackermann in 1928 [2].Entscheidungsproblem is about finding an effective procedure (algorithm) by which, given any expression $M$ in the notation of the system, it can be determined whether or not $M$ is provable in the system.Alan Turing and Alonzo Church, in 1930’s, independently proved that Entscheidungsproblem is unsolvable [3, 4].To capture the notion of effective procedures, Alan Turing and Alonzo Church developed Turing machines and $\lambda$-calculus, respectively.General recursive functions and Post machines are some other formal notions of effective procedures that were also presented in 1930’s [5, 6].Lambda calculus, Turing machines, and other notions of effective procedures are equivalent in their computational powers because they can simulate each other[7, 8].Only Turing machines are briefly described here because of their relevance to this work.
A Turing machine consists of a finite set of states (control unit), a read-write head, and an infinite tape.A simple schematic of single tape Turing machine is shown in Fig. 1. At each step, a Turing machine reads the alphabet from tape cell that is under read-write head, changes its state, writes an alphabet on the tape cell that is under read-write head, and moves the tape head to the left $(L)$ or to the right $(R)$.If $Q$ is finite set of states and $\Gamma$ represents finite set of tape alphabets, then we can write transition function ($\delta$) of single tape Turing machine in following way.
$\delta:Q\times\Gamma\longrightarrow Q\times\Gamma\times\{L,R\}$
Unit time is required to execute a complete transition of transition function $\delta$.Set of input alphabet is subset of tape alphabet $\Gamma$.In the multi-tape Turing machine each tape has separate read-write head.We can write transition function of $k$-tape Turing machine in following way.
$\delta:Q\times\Gamma^{k}\longrightarrow Q\times\Gamma^{k}\times\{L,R\}^{k}$
If a problem is computable then Turing machine halts in some finite time and outputs right answer for given input.We can compute some problems more efficiently if we increase the number of tapes or alphabet size (i.e., $|\Gamma|$) [9].A single tape Turing machine can simulate $k$-tape Turing machine with quadratic slow down.If a $k$-tape Turing machine can compute a problem in $T(n)$ time then there exists a single tape Turing machine that can compute same problem in time $T(n^{2})$.For a given problem, we can achieve linear speedup (through tape compression) by increasing number of tape alphabets.If there is a problem that can be computed by some Turing machine in time $f(n)$ then, for any constant $\epsilon>0$, there exists a Turing machine with more tape alphabets that can compute the problem in time $\epsilon f(n)+n+2$.These quadratic/constant factor speedups, that we can achieve by increasing number of tapes or alphabet size, will play central role in this work.
Figure 1: Single tape Turing machine
Some variants of Turing machines appear computationally more powerful than other variants; but all these variants (defined over finite set of states) have been proven equally powerful because they can simulate each other, efficiently.This work revisits and examines the computational characteristics of different variants of Turing machines at very fine-grain level.We achieve this goal in two steps. I.Constructing a transform technique for Turing computable problems that transforms computable problems into another type of problems. II.Presenting the transformed problem to different variants of Turing machines for computation and then analysing computation of those transformed problems.The scope of this work is being restricted and it includes only deterministic variants Turing machines.
2 Construction of Transform
Suppose there is a deterministic single tape Turing machine $D_{R}$ and it is defined over binary input.The unit time that $D_{R}$ takes to complete one transition of its transition function is represented by $T_{t}$, and the time $T_{t}$ is constant and same for every Turing machine mentioned in this paper.Turing machine $D_{R}$ computes a partially computable function $R$ that is not total.For a given input, the Turing machine $D_{R}$ requires $N_{R}$ number of steps to compute $R$.If $R$ is computable for given input then $N_{R}$ is some finite number, and it depends on input length.For a given input, the function $R$ can be either computable or uncomputable.Total time required to compute $R$ for a given input is represented by $f(R)$.
$f(R)=N_{R}T_{t}$ (1)
With a given input, there are two possible outcomes for function $R$, either $R$ is a computable function (if $D_{R}$ computes and halts) or $R$ is uncomputable function (if $D_{R}$ loops).We can state these two possibilities in term of $f(R)$ in following way.For a given input
  1. I.
    If $R$ is computable then $f(R)$ has finite value
  2. II.
    If $f(R)$ has infinite value (i.e., $N_{R}\rightarrow\infty$ ) then $R$ is uncomputable
Suppose, function $P$ in (2) is a computable (partially computable and total) function.If $P$ is computable then $f(P)$ has some finite value for every input.
$f(P)=N_{P}T_{t}$ (2)
Suppose, we have a variable $F$ such that $f(P,F)=N_{P}T_{t}F_{P}$ and $1\leq F_{P}<\infty$.If $M_{P}=\nicefrac{{1}}{{F_{P}}}$ then $f(P,M)=\nicefrac{{N_{P}T_{t}}}{{M_{P}}}$ and $0<M_{P}\leq 1$.This variable $M$ will be explained as we move forward in construction of this transform.The behaviour of (3), when $M_{P}=1$ and $M_{P}\rightarrow 0$, is represented in (4) and (5), respectively.
$f(P,M)=\frac{N_{P}T_{t}}{M_{P}}$ (3)
$\lim_{{M_{P}=1}}f(P,M)=f(P)$ (4)
$\lim_{{M_{P}\to 0}}f(P,M)\rightarrow\infty$ (5)
Suppose, the new function $f(P,M)$ represents time required to compute some partially computable function $P(m)$ for a given input.Given that $P$ is computable function and $P(m)$ is partially computable function, we can summarise (4 and 5) in following way.
For a given input
  1. I.
    If $M_{P}=1$ then $f(P,M)=f(P)$; consequently, $P(m)$ is a computable function
  2. II.
    If $M_{P}\rightarrow 0$ then $f(P,M)\rightarrow\infty$; consequently, $P(m)$ is an uncomputable function
In (3), if $M_{P}=1-2^{-X_{P}}$ and $0<M_{P}\leq 1$ then $X_{P}=-\log_{2}(1-W)$ and $0<X_{p}\leq\infty$. After replacing $M_{P}$ with $1-2^{-X_{P}}$ in (3) we get (6).Given that $P$ is computable function and $P(x)$ is partially computable function.
$f(P,X)=\frac{N_{P}T_{t}}{1-2^{-X_{P}}}$ (6)
$\lim_{{X_{P}=\infty}}f(P,X)=f(P)$ (7)
$\lim_{{X_{P}\to 0}}f(P,X)\to\infty$ (8)
If it is given that $P$ is computable function and $P(x)$ is partially computable function then we can summarise (7 and 8) in following way.
For a given input
  1. I.
    If $X_{p}=\infty$ then $f(P,X)=f(P)$; consequently, $P(x)$ is a computable function
  2. II.
    If $X_{p}\rightarrow 0$ then $f(P,X)\to\infty$; consequently, $P(x)$ is an uncomputable function
We can implement partially computable and computable (partially computable and total) functions through Turing machines.For a given input, there always exists a Turing machine $M_{1}$ that can execute (a function or a decision) task $Y$ if there exists a partial computable function for $Y$ and $M_{1}$ halts on the given input.Now we present Stipulation machine that is also a realizable computational object, just like Turing machines; and Turing machines interacts with Stipulation machines through some mutually binding set of rules (postulates).We can materialize the mathematical formulations presented in (6,7 and 8) through a realizable computational scheme that consists of Turing machine, Stipulation machine and some postulates.
2.1 Stipulation machine
There is a Turing machine $M_{1}$ that computes problem $Y$, and it is connected to Stipulation machine.A simple schematic of interconnection between Turing machine and Stipulation machine is shown in Fig.2. There are eight postulates that describe the interactions between Turing machine and Stipulation machine. I. Stipulation machine writes input on the space specified for input in Turing machine $M_{1}$, and there is one and only space for input in $M_{1}$. II. Stipulation machine writes only those inputs that $M_{1}$ can decide, i.e., $M_{1}$ is total Turing machine for the Stipulation machine. III.Stipulation machine rewrites input in Turing machine $M_{1}$ after every fixed time interval $T_{Y}$. IV.Length of every recurring input is equal to input length of first input of a computation. V. No two consecutive inputs of Turing machine $M_{1}$ can be identical in a computation. VI. When Turing machine $M_{1}$ halts, it informs Stipulation machine. VII. When Stipulation machine receives halting signal, it stops rewriting input on Turing machine $M_{1}$ for the computation. VIII. We can write input in $M_{1}$ only through Stipulation machine when it is connected to Stipulation machine.
The time $T_{Y}$ is characteristic of Stipulation machine and $0<T_{Y}\leq\infty$, and $T_{Y}=\infty$ means Stipulation machine does not rewrite input.If there is a Turing machine $M_{1}$ that is connected with Stipulation machine then it is represented by $M_{1}(t)$, and $\langle M_{1}(t)\rangle=\langle M_{1}\rangle$.Just like Turing machine, Stipulation machine is also a realizable object.Therefore, we can construct the computational scheme that is based on the eight postulates and interactions presented in Fig. 2, and we can observe the conclusions that are based this computational scheme.
Figure 2: Flow of interactions between Turing machine and Stipulation machine
2.2 Computation by a Turing machine $M(t)$
There is a Turing machine $P_{r}$ that accepts input if sum of input binary string is even and it rejects if sum of input binary strings is odd.Turing machine $P_{r}$ is a decider if input is $a$ for all $a\in A$.
$A=\{1101+101,11+10111,101+1101,1+11+1011,110+111+100,\ldots\}$
We give an input string $a$ (e.g., $11+10111$) to $P_{r}$.The Turing machine $P_{r}$ decides and halts when given input is $a$, i.e., $P_{r}(a)$ is decidable.Now, we transform $P_{r}$ into $P_{r}(t)$ by connected it to Stipulation machine and give input $a$ through Stipulation machine. In case of $P_{r}(t)$ with input $a$, the behaviour of the decidability is quite different when $T_{P_{r}}$ is relatively large and $T_{P_{r}}$ is relatively small.In following two cases we analyse the decidability of $P_{r}(t)$ with input $a$ and $0<T_{P_{r}}\leq\infty$; particularly, when $T_{P_{r}}\to 0$ and $T_{P_{r}}=\infty$.
Case I. Decidability of $P_{r}(t)$ with input $a$ and relatively large $T_{P_{r}}$.
Stipulation machine writes input string $a$ on Turing machine $P_{r}(t)$, the Turing machine $P_{r}(t)$ decides and halts on input $a$ in time less than $T_{P_{r}}$, Stipulation machines receives halts signal from $P_{r}(t)$ and does not rewrite input.If $T_{P_{r}}$ is relatively large then there is no difference between $P_{r}$ and $P_{r}(t)$, i.e., both $P_{r}(a)$ and $P_{r}(t)(a)$ are decidable.
$L_{1}=\{\langle P_{r}(t),a\rangle|\text{ Turing machine }P_{r}(t)\text{ decides }a\}$
With input $a$ and relatively large $T_{P_{r}}$ the language $L_{1}$ decidable; both, $P_{r}(a)$ and $P_{r}(t)(a)$ are decidable.There is no difference between $P_{r}$ and $P_{r}(t)$ when $T_{P_{r}}=\infty$.Therefore, $P_{r}(t)=P_{r}$ when $T_{P_{r}}=\infty$.
Case II. Decidability of $P_{r}(t)$ with input $a$ and $T_{P_{r}}\to 0$.
If $T_{P_{r}}$ is relatively too small then inputs of $P_{r}(t)$ change too quickly.Now, the time $T_{P_{r}}$ is relatively so small that Stipulation machine rewrites input on $P_{r}(t)$ far before $P_{r}(t)$ decides and halts on an input.Consequently, $P_{r}(t)$ carries out computation on some parts of each recurring input in some order.The transitions that $P_{r}(t)$ takes depends on its transition function, current configuration and those some parts of recurring inputs.Those some parts and some order make arbitrary string $f$ of some arbitrary length, and composition of $f$ depends on $T_{P_{r}}$.The string $f$ may not be element of set $A$.
$L_{2}=\{\langle P_{r}(t),f\rangle|\text{ Turing machine }P_{r}(t)\text{ decides }f\}$
The language $L_{2}$ is undecidable because decidability of a Turing machine on an arbitrary string is undecidable. So, the decidability of $P_{r}(t)(a)$ depends on $T_{P_{r}}$. If $T_{P_{r}}$ is relatively large then $P_{r}(t)(a)$ is decidable but if $T_{P_{r}}$ is too small (i.e., $T_{P_{r}}\to 0$) then $P_{r}(t)(a)$ is undecidable.
Only end points of $T_{P_{r}}$ (i.e., $T_{P_{r}}\to 0$ and $T_{P_{r}}=\infty$) are considered in following subsection but $T_{P_{r}}$ can have any value on interval $(0\quad\infty]$ and it is discussed in next section (Section 3).The time $P_{r}$ takes to recognize $a$ in $P_{r}(a)$ is represented by $f(P_{r})$, and the time $P_{r}(t)$ takes to recognize $a$ in $P_{r}(t)(a)$ is represented by $f(P_{r},T)$.Following mathematical expression can represent the conclusions of case I and II.
$f(P_{r},T)=\frac{N_{P_{r}}T_{t}}{1-2^{-T_{P_{r}}}}$ (9)
$\lim_{{T_{P_{r}}=\infty}}f(P_{r},T)=f(P_{r})$ (10)
$\lim_{{T_{P_{r}}\to 0}}f(P_{r},T)\to\infty$ (11)
Given that $P_{r}(a)$ is decidable
  1. I.
    If $T_{P_{r}}=\infty$ then $f(P_{r},T)=f(P_{r})$; consequently, $P_{r}(t)(a)$ is decidable
  2. II.
    If $T_{P_{r}}\rightarrow 0$ then $f(P_{r},T)\to\infty$; consequently, $P_{r}(t)(a)$ is undecidable
Stipulation machine and Turing machine are realizable computational objects.Therefore, Stipulation machines and Turing machines can be designed to replicate these phenomena and conclusions.Now we describe some examples where there is some problem $P$ that is computable for input $\psi$, and we transform $P(\psi)$ into $P(t)(\psi)$.The problem $P$ can be a decision or a function problem.
2.3 Examples of Transformed Problems
The problem $P_{1}$ is to twice evaluate the satisfiability of a given formula (e.g., $\Phi$ in 10) for a given input ($n_{\Phi}$), if both answers are yes then accept, if both answers are no then reject; and if both answers are no same then start evaluating again and continue until we have two consecutive same answers.If $P_{1}$ is decidable with given input ($\Phi$ and $n_{\Phi}$) then it will remain decidable if we evaluate the satisfiability for any finite number of times instead of two times, because we will have same answers (yes/no) irrespective how many time we evaluate given formula for the given input. We transform $P_{1}$ into $P_{1}(t)$, and variation time of Stipulation machine for $P_{1}(t)$ is represented by $T_{P_{1}}$.We give $\Phi$ and $n_{\Phi}$ for which the problem $P_{1}$ is decidable, and then through stipulation machine we give same input ($\Phi$ and $n_{\Phi}$) to $P_{1}(t)$
$\Phi=(x_{1}\lor\overline{x_{5}}\lor x_{7})\land(\overline{x_{2}}\lor\overline{x_{4}}\lor x_{8})\land(\overline{x_{2}}\lor\overline{x_{3}}\lor x_{6})$ (12)
$\Phi_{d}=(x_{1}(t)\lor\overline{x_{5}(t)}\lor x_{7}(t))\land(\overline{x_{2}(t)}\lor\overline{x_{4}(t)}\lor x_{8}(t))\land(\overline{x_{2}(t)}\lor\overline{x_{3}(t)}\lor x_{6}(t))$ (13)
$f(P_{1},T)=\frac{N_{P_{1}}T_{t}}{1-2^{-T_{P_{1}}}}$ (14)
$\lim_{{T_{P_{1}}=\infty}}f(P_{1},T)=f(P_{1})$ (15)
$\lim_{{T_{P_{1}}\to 0}}f(P_{1},T)\to\infty$ (16)
Given that ${P_{1}}(\Phi,n_{\Phi})$ is decidable
  1. I.
    If $T_{P_{1}}=\infty$ then $f(P_{1},T)=f(P_{1})$; consequently, $P_{1}(t)(\Phi,n_{\Phi})$ is decidable problem
  2. II.
    If $T_{P_{1}}\rightarrow 0$ then $f(P_{1},T)\to\infty$; consequently, $P_{1}(t)(\Phi,n_{\Phi})$ is undecidable problem
Through this transform technique, we can map a decidable decision problem into a decidable problem or an undecidable problem.Suppose there are three $n\times n$ non-zero matrices $A$, $B$, and $C$ such that $AB=C$.The problem $P_{2}$ is to compute $C$ four times for given $A$ and $B$, if in all four answers $(c_{11}\geq c_{nn})\land(c_{1n}<c_{n1})$ then accept, if in all four answers $(c_{11}<c_{nn})\land(c_{1n}\geq c_{n1})$ then reject; and if all four answers are not same then compute $C$ again from start and continue until we have four consecutive same answers.We transform $P_{2}$ into $P_{2}(t)$, and variation time of Stipulation machine for $P_{2}(t)$ is represented by $T_{P_{2}}$. We give $A$ and $B$ for which $P_{2}(A,B)$ is decidable, and then we gave same input $(A,B)$ to $P_{2}(t)$ through Stipulation machine.
{paracol}
2
$A=\begin{bmatrix}a_{11}&a_{12}&a_{13}&\cdots&a_{1n}\\ a_{21}&a_{22}&a_{23}&\cdots&a_{2n}\\ a_{31}&a_{32}&a_{33}&\cdots&a_{3n}\\ \vdots&\vdots&\vdots&\ddots&\vdots\\ a_{n1}&a_{n2}&a_{n3}&\cdots&a_{nn}\end{bmatrix}$
[…]
$B=\begin{bmatrix}b_{11}&b_{12}&b_{13}&\cdots&b_{1n}\\ b_{21}&b_{22}&b_{23}&\cdots&b_{2n}\\ b_{31}&b_{32}&b_{33}&\cdots&b_{3n}\\ \vdots&\vdots&\vdots&\ddots&\vdots\\ b_{n1}&b_{n2}&b_{n3}&\cdots&b_{nn}\end{bmatrix}$
$\begin{bmatrix}a_{11}&a_{12}&a_{13}&\cdots&a_{1n}\\ a_{21}&a_{22}&a_{23}&\cdots&a_{2n}\\ a_{31}&a_{32}&a_{33}&\cdots&a_{3n}\\ \vdots&\vdots&\vdots&\ddots&\vdots\\ a_{n1}&a_{n2}&a_{n3}&\cdots&a_{nn}\end{bmatrix}\times\begin{bmatrix}b_{11}&b_{12}&b_{13}&\cdots&b_{1n}\\ b_{21}&b_{22}&b_{23}&\cdots&b_{2n}\\ b_{31}&b_{32}&b_{33}&\cdots&b_{3n}\\ \vdots&\vdots&\vdots&\ddots&\vdots\\ b_{n1}&b_{n2}&b_{n3}&\cdots&b_{nn}\end{bmatrix}=\begin{bmatrix}c_{11}&c_{12}&c_{13}&\cdots&c_{1n}\\ c_{21}&c_{22}&c_{23}&\cdots&c_{2n}\\ c_{31}&c_{32}&c_{33}&\cdots&c_{3n}\\ \vdots&\vdots&\vdots&\ddots&\vdots\\ c_{n1}&c_{n2}&c_{n3}&\cdots&c_{nn}\end{bmatrix}$
$f(P_{2},T)=\frac{N_{P_{2}}T_{t}}{1-2^{-T_{P_{2}}}}$ (17)
$\lim_{{T_{P_{2}}=\infty}}f(P_{2},T)=f(P_{2})$ (18)
$\lim_{{T_{P_{2}}\to 0}}f(P_{2},T)\to\infty$ (19)
Given that $P_{2}(A,B)$ decidable
  1. I.
    If $T_{P_{2}}=\infty$ then $f(P_{2},T)=f(P_{2})$; consequently, $P_{2}(t)(A,B)$ is decidable
  2. II.
    If $T_{P_{3}}\to 0$ then $f(P_{3},T)\to\infty$; consequently, $P_{3}(t)(A,B)$ is undecidable
Suppose there are two $n\times n$ matrices $A$ and $B$, and both matrices have integer entries.The problem $P_{3}$ is to compute $AB$ thrice, and in all three computations, if $AB=I$ then accept, if $AB\neq I$ then reject; and if all three answers are not same then compute $AB$ again from start and continue until we have three consecutive same answers.We transform $P_{3}$ into $P_{3}(t)$, and variation time of Stipulation machine for $P_{3}(t)$ is represented by $T_{P_{3}}$. We give $A$ and $B$ for which $P_{3}(A,B)$ is decidable, and then we gave same input to $P_{3}(t)$ through Stipulation machine. {paracol}2
$A=\begin{bmatrix}a_{11}&a_{12}&a_{13}&\cdots&a_{1n}\\ a_{21}&a_{22}&a_{23}&\cdots&a_{2n}\\ a_{31}&a_{32}&a_{33}&\cdots&a_{3n}\\ \vdots&\vdots&\vdots&\ddots&\vdots\\ a_{n1}&a_{n2}&a_{n3}&\cdots&a_{nn}\end{bmatrix}$
[…]
$B=\begin{bmatrix}b_{11}&b_{12}&b_{13}&\cdots&b_{1n}\\ b_{21}&b_{22}&b_{23}&\cdots&b_{2n}\\ b_{31}&b_{32}&b_{33}&\cdots&b_{3n}\\ \vdots&\vdots&\vdots&\ddots&\vdots\\ b_{n1}&b_{n2}&b_{n3}&\cdots&b_{nn}\end{bmatrix}$
$\begin{bmatrix}a_{11}&a_{12}&a_{13}&\cdots&a_{1n}\\ a_{21}&a_{22}&a_{23}&\cdots&a_{2n}\\ a_{31}&a_{32}&a_{33}&\cdots&a_{3n}\\ \vdots&\vdots&\vdots&\ddots&\vdots\\ a_{n1}&a_{n2}&a_{n3}&\cdots&a_{nn}\end{bmatrix}\times\begin{bmatrix}b_{11}&b_{12}&b_{13}&\cdots&b_{1n}\\ b_{21}&b_{22}&b_{23}&\cdots&b_{2n}\\ b_{31}&b_{32}&b_{33}&\cdots&b_{3n}\\ \vdots&\vdots&\vdots&\ddots&\vdots\\ b_{n1}&b_{n2}&b_{n3}&\cdots&b_{nn}\end{bmatrix}=\begin{bmatrix}1&0&0&\cdots&0\\ 0&1&0&\cdots&0\\ 0&0&1&\cdots&0\\ \vdots&\vdots&\vdots&\ddots&\vdots\\ 0&0&0&\cdots&1\end{bmatrix}$
$f(P_{3},T)=\frac{N_{P_{3}}T_{t}}{1-2^{-T_{P_{3}}}}$ (20)
$\lim_{{T_{P_{3}}=\infty}}f(P_{3},T)=f(P_{3})$ (21)
$\lim_{{T_{P_{3}}\to 0}}f(P_{3},T)\to\infty$ (22)
Given that $P_{3}(A,B)$ is decidable
  1. I.
    If $T_{P_{3}}=\infty$ then $f(P_{3},T)=f(P_{3})$; consequently, $P_{3}(t)(A,B)$ is decidable
  2. II.
    If $T_{P_{2}}\to 0$ then $f(P_{2},T)\to\infty$; consequently, $P_{3}(t)(A,B)$ is undecidable
So far, we have applied this transform technique on decidable decision problems, but we can also apply it on computable functions problems.Suppose there is a set $u$ that consists of finite number of finite length binary strings, and each string has equal number of ones.The problem $P_{4}$ is to compute the number of ones in each string, if number of ones are equal in each string then output the result and halt, and if number of ones in each string are not equal then start computing again and continue until equal number of ones computed in each string.We transform $P_{4}$ into $P_{4}(t)$, and variation time of Stipulation machine for $P_{4}(t)$ is represented by $T_{P_{4}}$.The problem $P_{4}$ is computable for input $u$, and we give input $u$ to $P_{4}(t)$ through Stipulation machine.
$f(P_{4},T)=\frac{N_{P_{4}}T_{t}}{1-2^{-T_{P_{4}}}}$ (23)
$\lim_{{T_{P_{4}}=\infty}}f(P_{4},T)=f(P_{4})$ (24)
$\lim_{{T_{P_{4}}\to 0}}f(P_{4},T)\to\infty$ (25)
Given that $P_{4}(u)$ computable
  1. I.
    If $T_{P_{4}}=\infty$ then $f(P_{4},T)=f(P_{4})$; consequently, $P_{4}(t)(u)$ is computable
  2. II.
    If $T_{P_{4}}\to 0$ then $f(P_{4},T)\to\infty$; consequently, $P_{4}(t)(u)$ is uncomputable
Suppose there is an $m\times n$ matrix A with $a_{ij}\in\mathbb{N}$.The problem $P_{5}$ is to compute $det(A)$ through Gaussian elimination, and it is a computable problem when input is $A$.We transform $P_{5}$ into $P_{5}(t)$, and variation time of Stipulation machine for $P_{5}(t)$ is represented by $T_{P_{5}}$.We give $A$ for which $P_{5}$ is computable, and then we give $A$ to $P_{5}(t)$ through Stipulation machine.
$|A|=\left|\begin{array}[]{cccccc}a_{11}&a_{12}&a_{13}&\cdots&a_{1n}\\ a_{21}&a_{22}&a_{23}&\cdots&a_{2n}\\ a_{31}&a_{32}&a_{33}&\cdots&a_{3n}\\ \vdots&\vdots&\vdots&\ddots&\vdots\\ a_{m1}&a_{m2}&a_{m3}&\cdots&a_{mn}\end{array}\right|$ (26)
$|A_{d}|=\left|\begin{array}[]{cccccc}a_{11}(t)&a_{12}(t)&a_{13}(t)&\cdots&a_{1n}(t)\\ a_{21}(t)&a_{22}(t)&a_{23}(t)&\cdots&a_{2n}(t)\\ a_{31}(t)&a_{32}(t)&a_{33}(t)&\cdots&a_{3n}(t)\\ \vdots&\vdots&\vdots&\ddots&\vdots\\ a_{m1}(t)&a_{m2}(t)&a_{m3}(t)&\cdots&a_{mn}(t)\end{array}\right|$ (27)
Given that $P_{5}(A)$ computable
  1. I.
    If $T_{P_{5}}=\infty$ then $f(P_{5},T)=f(P_{5})$; consequently, $P_{5}(t)(A)$ is computable
  2. II.
    If $T_{P_{5}}\to 0$ then $f(P_{5},T)\to\infty$; consequently, $P_{5}(t)(A)$ is uncomputable
Suppose there is an $n\times n$ matrix $Q$ with $q_{ij}\in\mathbb{N}$, and rank of $Q$ is $n$.The problem $P_{5}$ is to compute inverse of a given matrix through Gaussian elimination, and $P_{5}$ is computable when input is $Q$.Because $Q$ is invertible matrix and $(Q^{-1})^{-1}=Q$ therefore $P_{6}(P_{6}(Q))=Q$. If $A$ and $B$ are two invertible matrices and $A\neq B$ then $P_{6}(A)\neq P_{6}(B)$.We transform $P_{6}$ into $P_{6}(t)$, and variation time of Stipulation machine for $P_{6}(t)$ is represented by $T_{P_{6}}$.We give $Q$ for which $P_{6}$ is computable, and then we give $Q$ to $P_{6}(t)$ through Stipulation machine.
$Q=\begin{bmatrix}q_{11}&q_{12}&q_{13}&\cdots&q_{1n}\\ q_{21}&q_{22}&q_{23}&\cdots&q_{2n}\\ q_{31}&q_{32}&q_{33}&\cdots&q_{3n}\\ \vdots&\vdots&\vdots&\ddots&\vdots\\ q_{n1}&q_{n2}&q_{n3}&\cdots&q_{nn}\end{bmatrix}$
Given that $P_{6}(Q)$ computable
  1. I.
    If $T_{P_{6}}=\infty$ then $f(P_{6},T)=f(P_{6})$; consequently, $P_{6}(t)(Q)$ is computable
  2. II.
    If $T_{P_{6}}\to 0$ then $f(P_{6},T)\to\infty$; consequently, $P_{6}(t)(Q)$ is uncomputable
Suppose there is a problem that has partially computable function $P$, and $P$ is computable when input is $\psi$. We transform $P(\psi)$ into $P(t)(\psi)$ through a transform that will be represented by $f(Z,T)$; and transformed problem $P(t)(\psi)$ can be computable or uncomputable, depending upon time $T_{P}$.
$P(\psi)$ (Computable)$f(Z,T)\hskip 5.0ptTransform$$P(t)(\psi)$ (Computable if $T_{P}=\infty$)$P(t)(\psi)$ (Uncomputable if $T_{P}\to 0$)
The time $T_{P}\in(0\text{ }\infty]$ but we have discussed only two cases, i.e., $T_{P}=\infty$ and $T_{P}\to 0$. Stipulation machine can have any value of $T_{P}$ on interval $(0\text{ }\infty]$.Now we discuss the behaviour of $f(Z,T)$ transform when $T_{P}$ can have any possible value on interval $(0\text{ }\infty]$ for transformed problem $P(t)(\psi)$.
$P(\psi)$ (Computable)$f(Z,T)\hskip 5.0ptTransform$$P(t)(\psi)$ (Computable if $T_{P}=\infty$)$P(t)(\psi)$ (Uncomputable if $T_{P}\to 0$)$P(t)(\psi)$ (If $T_{P}\in(0\quad\infty)$ then?)
3 Normalization of $f(Z,T)$ Transform
Suppose there is problem $P$ that is computable for input $\psi$, and $\psi\in\Psi$. We apply $f(Z,T)$ transform on $P(\psi)$ and get $P(t)(\psi)$.Normalization of the transformed problem $P(t)(\psi)$ is defined in following way.
$f(P,T)=\frac{N_{P}T_{t}}{1-2^{-T_{P}}}$ (28)
$f_{n}(P,T)=\frac{f(P,T)}{\frac{T_{P}}{1-2^{-T_{P}}}}$ (29)
$f_{n}(P,T)=\frac{N_{P}T_{t}}{T_{P}}$ (30)
Because $P(\psi)$ is computable problem and $0<T_{P}\leq\infty$ therefore $0\leq f_{n}(P,T)<\infty$.The problem $P(t)(\psi)$ is computable if $10\leq f_{n}(P,T)<$ 1 because $P(t)(\psi)$ halts before Stipulation machine rewrites input. If $N_{P}T_{t}>T_{P}$ then $f_{n}(P,T)>1$ i.e., input of $P(t)$ varies before $P(t)(\psi)$ gets computed and $P(t)$ halts. Given that $\psi\in\Psi$ and $\Psi=\{\psi_{1},\text{ }\psi_{2},\text{ }\psi_{3},\text{ }\ldots,\text{ }\psi_{n}\}$, and according to postulate V, no two consecutive inputs of $P(t)$ are identical.If $f_{n}(P,T)>1$ then Stipulation machine rewrites input on $P(t)$ before $P(t)$ decides and halts on input $\psi$.Consequently, $P(t)$ carries out computation on some parts of each recurring input in some order.Those some parts and some order make arbitrary string $\psi_{f}$ of some arbitrary length, and composition of $\psi_{f}$ depends on $T_{P}$.The string $\psi_{f}$ may not be member of set $\Psi$.Therefore, if $f_{n}(P,T)>1$ then $P(t)$ is not computable when input is $(\psi)$, but if $10\leq f_{n}(P,T)<$ 1 then $P(t)$ is computable when input is $(\psi)$.Normalization line for $P(t)(\psi)$ is shown in Fig.3.The set $\Psi$ can have input strings of any length, and this important aspect (infinite cardinality of set $\Psi$) is discussed in Section 5.
Figure 3: Normalization line of $f(Z,T)$ Transform
Point of inflection and point of convergence are two important points on normalization line. If $f_{n}(P,T)=0$ then transformed problem $P(t)(\psi)$ exist at point of convergence, and $f(P,T)=f(P)$ when $f_{n}(P,T)=0$.If $f_{n}(P,T)=1$ then transformed problem $P(t)(\psi)$ exist at point of inflection.The point where $f_{n}(P,T)=1$ is very critical point for computability of problem $P(t)(\psi)$.A small change can make $P(t)(\psi)$ move in-between computable-uncomputable when $f_{n}(P,T)$ exists around point of inflection.If $P(\psi)$ requires $N_{P}$ steps and $P(t)(\psi)$ exists around point of inflection, then slight change in $N_{P}$ (e.g., $N_{P}-n$ or $N_{P}+n$ for some $n\geq 1$) or $T_{P}$ can change the computability of $P(t)(\psi)$.We can also represent normalization line of $P(t)(\psi)$ in following way.
$P(\psi)$ (Computable)$f(Z,T)\hskip 5.0ptTransform$$P(t)(\psi)$ (If $T_{P}=\infty$ then $f_{n}(P,T)=0$)$P(t)(\psi)$ ($10<f_{n}(P,T)<$ 1)$P(t)(\psi)$ (If $T_{P}\to 0$ then $f_{n}(P,T)\to\infty$)$P(t)(\psi)$ ($f_{n}(P,T)>1$)$P(t)(\psi)$ ($f_{n}(P,T)=1$ for some $T_{P}$)
The $f(Z,T)$ transformed problems that exist just before or immediately after point of inflection are very helpful in analysing the computational characteristics of different variants of the Turing machines at infinitesimal scale.If cardinality of $f(Z,T)$ transformed problems is higher than cardinality of Turing machines (defined over finite set of states) then there must exist $f(Z,T)$ transformed problems that are uncomputable.Cardinality of $f(Z,T)$ transformed problems is discussed in Section 5.
4 Computation of $f(Z,T)$ Transformed Problems by variants of Turing machines
[…]
Figure 4: Computability of $P(t)(\psi)$ through variants of Turing machines
4.1 $k$-tape Turing machines for $P(t)(\psi)$ problems
If there is a string $m$ of length $n$ then $1$-tape Turing machine requires $O(n^{2})$ number of steps to decide whether $m$ is palindrome or not, and $2$-tape Turing machine requires $O(n)$ steps to make this decision about $m$.The problem $P$ is to decide twice whether the given input string is palindrome, if both answers are yes then accept, if both answers are no then reject; and if both answers are not same then start evaluating again and continue until we have two consecutive same answers.For a given input $m$, we transform $P(m)$ into $P(t)(m)$, and then adjust $T_{P}$ in such a way that $f_{n}(P,T)<1$ for $2$-tape Turing machines and $f_{n}(P,T)$ of $P(t)(m)$ exists just before point of inflection.Single tape and $2$-tape Turing machines are available, and both can recognize $P(m)$.Now, for this $P(t)(m)$ and $f_{n}(P,T)$, replace $2$-tape Turing machine connected to Stipulation machine with $1$-tape Turing machine.The problem $P(t)(m)$ is decidable for $2$-tape Turing machine because $f_{n}(P,T)<1$, but it is not decidable by $1$-tape Turing machine because $N_{P}$ is higher for $1$-tape Turing machine (as it requires $O(n^{2})$ steps) and consequently $f_{n}(P,T)>1$.The transformed problem $P(t)(m)$ moves from being decidable (i.e., $f_{n}(P,T)<1$) for $2$-tape Turing machine to undecidable (i.e., $f_{n}(P,T)>1$) for $1$-tape Turing machine.If there is a $k$-tape Turing machine that can decide $P(m)$ then, only for $k\geq 2$, it can decide this $P(t)(m)$ with given $f_{n}(P,T)$.
Suppose there is a set $\Pi$ that consists of $k$ number of binary strings, and each string has some finite length $n$ for some $n\geq 10$.The problem $P_{1}$ is to decide thrice that the binary strings of $\Pi$ are identical, if all three answers are yes then accept, if all three answers are no then reject; and if all answers are not same then start evaluating again and continue until we have three consecutive same answers.Suppose $P_{1}$ is decidable for a given $\Pi$.We transform $P_{1}(\Pi)$ into $P_{1}(t)(\Pi)$, and then adjust $T_{P_{1}}$ in such a way that $f_{n}(P_{1},T)<1$ for $k$-tape Turing machines, and $f_{n}(P_{1},T)$ exists just before point of inflection for $P_{1}(t)(\Pi)$.Now, for this $P_{1}(t)(\Pi)$ and $f_{n}(P_{1},T)$, replace $k$-tape Turing machine connected to Stipulation machine with $k-1$ tape Turing machine. The problem $P_{1}(t)(\Pi)$ is decidable by $k$-tape Turing machine but it is not decidable by $k-1$ tape Turing machine because $N_{P_{1}}$ will increase for $k-1$ tape Turing machine and consequently $f_{n}(P_{1},T)>1$.The problem $P_{1}(t)(\Pi)$ moves from being decidable (i.e., $f_{n}(P_{1},T)<1$) for $k$-tape Turing machine to undecidable (i.e., $f_{n}(P_{1},T)>1$) for $k-1$ tape Turing machine.Similarly, there exists a transformed problem that is computable for $k+1$ tape Turing machine but not computable for $k$ tape Turing machine.
We can construct a computable problem $P$ that, for a given input $\psi$, takes slightly more number of steps in its computation if we decrease the number of tapes.This slight increment in the number of computational steps provides opportunity to apply $f(Z,T)$ transform and get transformed problem $P(t)(\psi)$, we adjust its variation time $T_{P}$ such that $f_{n}(P,T)$ exist around point of inflection for $P(t)(\psi)$, and then present $P(t)(\psi)$ to different variants of Turing machines for computation.Depending on the variant, the problem $P(t)(\psi)$ can be computable or uncomputable, because $f_{n}(P,T)$ for $P(t)(\psi)$ is different for different variants.
If $T(k,2)$ represents $k$-tape Turing machine defined over binary input, then there exist problem $P(t)(\psi)$ such that Turing machines can be enlisted according to their computational power in following way for $f(Z,T)$ transformed problems:
$T(1,2)\subsetneq T(2,2)\subsetneq T(3,2)\subsetneq\ldots\subsetneq T(k,2)\subsetneq T(k+1,2)\ldots$ (31)
We can state that there exists a problem $P(t)(\psi)$ such that $T(k,2)\neq T(k+1,2)$ because $f_{n}(P,T)>1$ for $T(k,2)$ and $f_{n}(P,T)<1$ for $T(k+1,2)$
For a given problem $P(t)(\psi)$, if $f_{n}(P,T)=0$ because $T_{P}=\infty$ then $T(1,2)=T(k,2)$ for all finite $k$, i.e., if input of problem $P(t)$ does not vary at all (because $T_{P}=\infty$) then single tape Turing machine is computationally as powerful as any other Turing machine defined over finite number of tapes.To put it in another way, if problem $P(t)(\psi)$ exist at point of convergence then all Turing machines (defined over finite number of tapes) converge to equal computational power.
4.2 $w$-alphabet Turing machines for $P(t)(\psi)$ problems
As stated earlier in linear speedup theorem that increase in alphabet size results in tape compression, and that consequently decreases the number of steps required in a computation.Any change in number of steps required in a computation changes the normalization value of $f(Z,T)$ transformed problem, and a change in the normalization value of a transformed problem can cause outcome of computability for the transformed problem to change.
We need unary string of length $2^{x}$ to represent binary string of length $x$.Turing machines defined over unary input can have better time complexity only if the time complexity is measured in term of unary input length, but if time complexity is measured in term of total number of steps then Turing machines defined over unary input take more number of steps to compute a problem than that of Turing machines defined over binary or higher numeral systems.
The problem $P$ is to decide whether the given input string $m$ is palindrome or not, if both answers are yes then accept, if both answers are no then reject; and if both answers are not same then start evaluating again and continue until we have two consecutive same answers. For given input string $m$, we transform this problem $P(m)$ into $P(t)(m)$, and then adjust $T_{P}$ in such a way that $f_{n}(P,T)<1$ for a Turing machine with input alphabet $w=\{0,1\}$, and $f_{n}(P,T)$ exists just before point of inflection for $P(t)(m)$. If $N_{P}$ represents number of steps required to compute $P(m)$, then decrease in $|w|$ will increase $N_{P}$ and normalization value. There exist $m$ such that $P(t)(m)$ is computable if $|w|>1$ because $f_{n}(P,T)<1$, and $P(t)(m)$ not computable if $|w|=1$ because $f_{n}(P,T)>1$. Similarly, we can construct and show the existence of $f(Z,T)$ transformed problem that is computable if $|w|\geq g+1$ and it is not computable if $|w|=g$ for any positive integer $g$.
We can construct a computable problem $P$ that, for a given input $\psi$, takes slightly fewer number of steps in its computation if we increase alphabet size. This slight decrement in the number of computational steps provides opportunity to apply $f(Z,T)$ transform and get transformed problem $P(t)(\psi)$, we adjust its variation time $T_{P}$ such that $f_{n}(P,T)$ exist around point of inflection for $P(t)(\psi)$, and then present $P(t)(\psi)$ to different variants of Turing machines for computation.Depending on the variant of Turing machine, the problem $P(t)(\psi)$ can be computable or uncomputable, because $f_{n}(P,T)$ for $P(t)(\psi)$ is different for different variants.
If $T(1,w)$ represents single tape $|w|$-input alphabet Turing machine, then there exist problem $P(t)(\psi)$ such that Turing machines can be enlisted according to their computational power in following way for $f(Z,T)$ transformed problems:
$T(1,1)\subsetneq T(1,2)\subsetneq T(1,3)\subsetneq\ldots\subsetneq T(1,w)\subsetneq T(1,w+1)\ldots$ (32)
We can state that there exists a problem $P(t)(\psi)$ such that $T(1,w)\neq T(1,w+1)$ because $f_{n}(P,T)>1$ for $T(1,w)$ and $f_{n}(P,T)<1$ for $T(1,w+1)$.
For a given problem $P(t)(\psi)$, if $f_{n}(P,T)=0$ (because $T_{P}=\infty$) then $T(1,1)=T(1,w)$ for all finite set of $w$. To describe it in another way, if problem $P(t)(\psi)$ exist at point of convergence then all Turing machines, defined over finite set of input alphabet, converge to equal computational power.
4.3 $k$-tape and $w$-alphabet Turing machines for $P(t)(\psi)$ problems
A $k$-tape Turing machine defined over $w$ input alphabet is represented by $T(k,w)$. As mentioned in subsection 4.1 and 4.2, we can construct and show the existence of $f(Z,T)$ transformed problem $P(t)(\psi)$ such that $f_{n}(P,T)>1$ for $T(n,n)$ and $P(t)(\psi)$ is not computable, but the problem $P(t)(\psi)$ has $f_{n}(P,T)<1$ for $T(n+1,n+1)$ and $P(t)(\psi)$ is computable.
We can construct a computable problem $P$ that, for a given input $\psi$, takes slightly fewer number of steps in its computation if we increase alphabet size and number of tapes.This slight decrement in the number of computational steps provides opportunity to apply $f(Z,T)$ transform and get problem $P(t)(\psi)$, we adjust variation time $T_{P}$ in such a way that $f_{n}(P,T)$ exist around point of inflection for $P(t)(\psi)$, and present $P(t)(\psi)$ to different variants of Turing machines for computation.Depending on the variant, the problem $P(t)(\psi)$ can be computable or uncomputable, because $f_{n}(P,T)$ for $P(t)(\psi)$ is different for different variants.
There exist problem $P(t)(\psi)$ such that Turing machines $T(k,w)$ can be enlisted according to their computational power in following way for $f(Z,T)$ transformed problems:
$T(1,1)\subsetneq T(2,2)\subsetneq T(3,3)\subsetneq\ldots\subsetneq T(n,n)\subsetneq T(n+1,n+1)\ldots$ (33)
We can state that there exists a problem $P(t)(\psi)$ such that $T(n,n)\neq T(n+1,n+1)$ because $f_{n}(P,T)>1$ for $T(n,n)$ and $f_{n}(P,T)<1$ for $T(n+1,n+1)$
For a given problem $P(t)(\psi)$, if $f_{n}(P,T)=0$ (because $T_{P}=\infty$) then $T(1,1)=T(n,n)$ for all finite $n$, i.e., if input of problem $P(t)$ does not change at all (because $T_{P}=\infty$) then Turing machine defined over $k=1$ and $w=1$ is computationally as powerful as any Turing machine defined over finite number of tapes and finite set of input alphabets.
It can be stated that if problem $P(t)(\psi)$ exists at point of convergence then all Turing machines (defined over finite number of tapes and finite set of input alphabet) converge to equal computational power.All variants of Turing machine have equal computational power if $f(Z,T)$ transformed problems exist at point of convergence, i.e., $f_{n}(Z,T)=0$.
If there exist $f(Z,T)$ transformed problems that are uncomputable by Turing machines then cardinality of $f(Z,T)$ transformed problems must be higher than cardinality of Turing machines.
5 Cardinality of Problems Produced by $f(Z,T)$ Transform
The $f(Z,T)$ transform is applied to only computable functions (or decidable languages) and it produces transformed problems.Suppose there is a decidable language $L_{C}$ and it has only finite number of strings, and each string in $L_{C}$ has length $n$, i.e., $l_{C}\in L_{C}$ and $|l_{C}|=n$ for all $l_{C}$.There exist a Turing machine $P_{C}$ that can recognize $L_{C}$, and $P_{C}(l_{C})$ is decidable for all $l_{C}$.We want to know the cardinality of $f(Z,T)$ transformed problems if we apply $f(Z,T)$ transform on $L_{C}$.We change $P_{C}$ into $P_{C}(t)$ and input is from $L_{C}$. If $0<T_{P_{C}}\leq\infty$ and $T_{P_{C}}\in\mathbb{Q}^{+}$ then $|T_{P_{C}}|=\aleph_{0}$.If $|T_{P_{C}}|=\aleph_{0}$ then $f(Z,T)$ transform produces infinite number of transformed problems in the form of $P_{C}(t)(l_{C})$ for $T_{P_{C}}(1)$, $P_{C}(t)(l_{C})$ for $T_{P_{C}}(2)$, $P_{C}(t)(l_{C})$ for $T_{P_{C}}(3)$, …$P_{C}(t)(l_{C})$ for $T_{P_{C}}(\infty)$, and $T_{P_{C}}=\{T_{P_{C}}(1),\text{ }T_{P_{C}}(2),\text{ }T_{P_{C}}(3),\text{ }\ldots,\text{ }T_{P_{C}}(\infty)\}$.The cardinality of $f(Z,T)$ transformed problems is countably infinite when we apply $f(Z,T)$ transform on a language that has only finite number of strings in it, e.g., language $L_{C}$.
$L_{C}=\{l_{C(1)},\text{ }l_{C(2)},\text{ }l_{C(3)},\text{ }l_{C(4)},\text{ }l_{C(5)},\text{ }\ldots,\text{ }l_{C(128)}\}$
$P_{C}(l_{C})$ (Decidable)$f(Z,T)\hskip 5.0ptTransform$$P_{C}(t)(l_{C})$ for $T_{P_{C}}(1)$$P_{C}(t)(l_{C})$ for $T_{P_{C}}(2)$$P_{C}(t)(l_{C})$ for $T_{P_{C}}(3)$$P_{C}(t)(l_{C})$ for $T_{P_{C}}(\infty)$
There is an other decidable language $L_{D}$ and cardinality of set that represents lengths of strings in $L_{D}$ is $\aleph_{0}$, and there exists a Turing machine $P_{D}$ that can recognize $L_{D}$.If $0<T_{P_{D}}\leq\infty$ and $T_{P_{D}}\in\mathbb{Q}^{+}$ then $|T_{P_{D}}|=\aleph_{0}$.We change $P_{D}$ into $P_{D}(t)$ and input of $P_{D}(t)$ is from $L_{D}$.If $|T_{P_{D}}|=\aleph_{0}$ then $f(Z,T)$ transform produces $\aleph_{0}^{\aleph_{0}}$ number of combinations, and each combination is $f(Z,T)$ transformed problem.There are uncountable number of $f(Z,T)$ transformed problems because $2^{\aleph_{0}}\leq\aleph_{0}^{\aleph_{0}}$.
$L_{D}=\{l_{D(1)},\text{ }l_{D(2)},\text{ }l_{D(3)},\text{ }l_{D(4)},\text{ }l_{D(5)}\text{ }\ldots\text{ }\}$
Suppose there is a decidable language $L_{E}$, and we apply $f(Z,T)$ transform on $L_{E}$.If $0<T_{E}\leq\infty$ and $n$ represents the length of strings in $L_{E}$ then Table $1$.represents $f(Z,T)$ transformed problems for language $L_{E}$.If $P_{E}$ is recognizer of $L_{E}$ and $l_{E}\in L_{E}$ then $P_{E}(t)(l_{E})=P_{E}(l_{E})$ for $T_{E}=\infty$ and $L(n\text{ }\infty)=L_{E}$.Therefore, it can be stated that the Language $L_{E}$ is subset of $f(Z,T)$ transformed $L_{E}$, and left most column of of following table represents language $L_{E}$.
Table 1: $f(Z,T)$ Transformed Problems for Language $L_{E}$
$\begin{array}[]{|c|c|c|c|c|c|c|c|c|c|}\hline\cr\mathrm{n}\backslash\mathrm{T_{E}}&1&2&3&4&5&6&7&\ldots&\infty\\ \hline\cr 1&L(1,1)&L(1,2)&L(1,3)&L(1,4)&L(1,5)&L(1,6)&L(1,7)&\ldots&L(1,\infty)\\ \hline\cr 2&{L}(2,1)&L(2,2)&L(2,3)&L(2,4)&L(2,5)&L(2,6)&L(2,7)&\ldots&L(2,\infty)\\ \hline\cr 3&L(3,1)&L(3,2)&L(3,3)&L(3,4)&L(3,5)&L(3,6)&L(3,7)&\ldots&L(3,\infty)\\ \hline\cr 4&L(4,1)&L(4,2)&L(4,3)&L(4,4)&L(4,5)&L(4,6)&L(4,7)&\ldots&L(4,\infty)\\ \hline\cr 5&L(5,1)&L(5,2)&L(5,3)&L(5,4)&L(5,5)&L(5,6)&L(5,7)&\ldots&L(5,\infty)\\ \hline\cr 6&L(6,1)&L(6,2)&L(6,3)&L(6,4)&L(6,5)&L(6,6)&L(6,7)&\ldots&L(6,\infty)\\ \hline\cr 7&L(7,1)&L(7,2)&L(7,3)&L(7,4)&L(7,5)&L(7,6)&L(7,7)&\ldots&L(7,\infty)\\ \hline\cr\vdots&\vdots&\vdots&\vdots&\vdots&\vdots&\vdots&\vdots&\ddots&\vdots\\ \hline\cr\infty&L(\infty,1)&L(\infty,2)&L(\infty,3)&L(\infty,4)&L(\infty,5)&L(\infty,6)&L(\infty,7)&\ldots&L(\infty,\infty)\\ \hline\cr\end{array}$
Cardinality of Turing machines defined over finite set of states is countably infinite and cardinality of $f(Z,T)$ transformed problems is uncountable, and this difference of cardinalities implies the existence of $f(Z,T)$ transformed problems that are uncomputable by Turing machines defined over finite set of states ( i.e., finite number of tapes and finite set of alphabet).
Through $f(Z,T)$ transform, we can demonstrate the existence of problems that can force different variants of Turing machines (that are defined over finite set of states) to exhibit, at infinitesimal scale, their differential computational characteristics.This work is strictly limited to deterministic variants of Turing machines and computability of $f(Z,T)$ transformed problems; and in future, it will be extended towards nondeterministic Turing machines and complexity theory.
6 Mathematical Formulation of Stipulation Machine Interactions
During construction of $f(Z,T)$ transform we assumed that $W_{P}=1-2^{-T_{P}}$ in (3). But if we assume that $W_{P}=1-2^{-\nicefrac{{T_{P}}}{{c}}}$ and $c$ is some positive constant then we get (34 and 35) after replacing $1-2^{-T_{P}}$ with $1-2^{-\nicefrac{{T_{P}}}{{c}}}$ in (6 and 29).
$f(P,T)=\frac{N_{P}T_{t}}{1-2^{-\nicefrac{{T_{P}}}{{c}}}}$ (34)
$f_{n}(P,T)=\frac{N_{P}T_{t}}{T_{P}}$ (35)
If constant $c=N_{P}T_{t}$ and we replace the value of $c$ with $N_{P}T_{t}$ in (34) then we get (36).We can describe computability of every $f(Z,T)$ Transformed problem through (35 and 36).There are four possible outcomes when we try to compute any $f(Z,T)$ transformed problems by a Turing machine (or by a Turing complete model of computation).
$f(P,T)=\frac{N_{P}T_{t}}{1-2^{-\left(\nicefrac{{T_{P}}}{{N_{P}T_{t}}}\right)}}$ (36)
[…]
Stipulation machine and Turing machine are realizable objects, and if they interact with each other through eight mutually binding rules (postulates) then their interactions can be described through $f(Z,T)$ transform.We apply $f(Z,T)$ transform only on computable problems, and transformed problems can be computable (case I and II) or uncomputable (case IV), and sometime the computability of the transformed problem is highly sensitive toward choice of computational scheme(case III).
7 Discussion
This paper presents a new realizable computational scheme and through this scheme we transform Turing computable problems into new type of problems.The transformed problems provide a window to analyse the computational characteristics of different variants of Turing machine at infinitesimal scale, and to enlist those variants of Turing machine according to their computational characteristics.This transform has some applied aspects as well.The $f(Z,T)$ transform describes interactions between Turing machine (or Turing complete models of computation) and Stipulation machine that take place under eight postulates.
There is a problem $P$ and it is computable on input $\psi$.We construct $f(Z,T)$ transform that transforms computable problem $P(\psi)$ into new problem $P(t)(\psi)$.Then we analyse the computability of transformed problem $P(t)(\psi)$ through different variants of Turing machines using $f_{n}(Z,T)$ normalization line.If we turn $P(\psi)$ into $P(t)(\psi)$ and $10\leq f_{n}(P,T)<$ 1 then $P(t)(\psi)$ is also computable just like $P(t)(\psi)$, but if $f_{n}(P,T)>1$ then $P(t)(\psi)$ is not computable problem.If $f_{n}(P,T)=1$ then $P(t)(\psi)$ exist at point of inflection, and this point provides a very important opportunity to investigate computational characteristics of different variants of Turing machines at infinitesimal scale.If we turn $P(\psi)$ into $P(t)(\psi)$ and $f_{n}(P,T)=0$ then $P(t)(\psi)$ exists at point of convergence.If $P(t)(\psi)$ exists at point of convergence then $T(k,w)=T(1,1)$ for every finite $k$ and $w$; i.e., finite number of additional tapes or additional alphabets do not provide additional computational power to a Turing machine.We apply $f(Z,T)$ transform on a computable problem $P(\psi)$ and we get transformed problem $P(t)(\psi)$.Following two equations represent the characteristics of this $f(Z,T)$ transformed problem.
$f(P,T)=\frac{N_{P}T_{t}}{1-2^{-\left(\nicefrac{{T_{P}}}{{N_{P}T_{t}}}\right)}}$ (37)
$f_{n}(P,T)=\frac{N_{P}T_{t}}{T_{P}}$ (38)
There are two expressions in (37 and 38) that represent time, i.e., $T_{t}$ and $T_{P}$.The time $T_{t}$ is characteristic of Turing machine and time $T_{P}$ is characteristic of Stipulation machine.If $(\nicefrac{{T_{t}}}{{T_{P}}})$ such that $f_{n}(P,T)>1$ then $P(t)(\psi)$ is uncomputable, and if $(\nicefrac{{T_{t}}}{{T_{P}}})$ such that $f_{n}(P,T)<1$ then $P(t)(\psi)$ is also computable, like $P(\psi)$. As stated earlier, if $f_{n}(P,T)=0$ then finite number of additional tapes or additional alphabet do not provide any additional computational power to Turing machines, and $P(t)(\psi)$ is computable by every Turing machine defined over finite number of tapes and finite set of alphabet. Therefore, it can be stated that established results in computability theory are special case of $f(Z,T)$ transform where $f_{n}(P,T)=0$. If problem $P(\psi)$ is a computable then problem $P(t)(\psi)$ with $T_{P}=\infty$ is also computable.Therefore, it can also be stated that $P(\psi)=P(t)(\psi)$ when $T_{P}=\infty$, and consequently, $T(1,1)=T(k,w)$, over finite number of $k$ and $w$.We can say that every $P(\psi)$ is in fact $P(t)(\psi)$ whose input never changes while it is being computed.The $f(Z,T)$ transform, through (37 and 38), establishes a notion of relativistic computability within computable problems.In this relative computability, if a problem is computable for a given inputs and inputs are changing with time, if the rate of change of inputs is sufficiently high for given model of computation, then the computable problem turns into uncomputable problem for the given model of computation. Suppose, there is a task $G$ and it is computable if and only if all given information (input, instructions, etc.) are taken into account.The task $G$ gets its information $\eta$ through an interface.We apply $f(Z,T)$ transform on $G(\eta)$ and we get (39 and 40).
$f(G,T)=\frac{N_{G}T_{t}}{1-2^{-\left(\nicefrac{{T_{G}}}{{N_{G}T_{t}}}\right)}}$ (39)
$f_{n}(G,T)=\frac{N_{G}T_{t}}{T_{G}}$ (40)
The $T_{G}$ in (39 and 40) represents time after which input information $\eta$ varies.The task $G(t)(\eta)$ with fixed input information has $f_{n}(G,T)=0$ because $T_{G}=\infty$ and it is computable task.If we start changing the input and $T_{t}\ll T_{G}$ such that $f_{n}(G,T)<1$ then $G(t)(\eta)$ remains a computable task, and if we start changing input in a way that $f_{n}(G,T)>1$ then $G(t)(\eta)$ turns into uncomputable task.If the task $G(t)(\eta)$ represents a fixed size phenomenon that evolves with time (as in 13 and 27 ) then $G(t)(\eta)$ is computable only if $f_{n}(G,T)<1$.Stipulation and Turing machines can be designed to replicate the phenomena that are based on $f(Z,T)$ transform.The phenomena that are based on $f(Z,T)$ Transform are realizable, and consequences of those phenomena are observable. Future work will extend the results of $f(Z,T)$ Transform toward complexity theory, and it will also present the applied aspects of this Transform.